An investor invested a total of 2400 in two mutual funds. One fund earned a 9% profit while the other earned 4% profit. If the investor's total was $161, how much was invested in each mutual fund?

Answer by checkley79(3341)

.09x+.04(2400-x)=161
.09x+96-.04x=161
.05x=161-96
.05x=65
x=65/.05
x=$1,300 was invested @ 9%.
2400-1300=$1100 invested @ 4%.
Proof:
.09*1300+.04*1100=161
117+44=161
161=161